OpenStudy (anonymous):

when give a graph of a quadratic function the questions asks you for zeros what are zeros

OpenStudy (moongazer):

zeros is the value of x when the equation is equal to zero

OpenStudy (moongazer):

for example what are the zeros of x^2+4x+4=0 (x+1)(x+3)=0 x=-1 x=-3

OpenStudy (moongazer):

click "good answer" if you are contented

OpenStudy (anonymous):

One other word is "roots", where the graph cross the x axis, and where the equation take the value 0.

OpenStudy (anonymous):

If the graph of a quadratic function is given, it will be of parabolic shape zeros are the points at which it intersects the x-axis

OpenStudy (anonymous):

so is it the point where it intersects or zeros is the value of x when the equation is equal to zero

OpenStudy (anonymous):

zero of a quadratic equation is the value(s) of x, these r given by the points where graph cuts x -axis .... see the attached file. in this the two red points on the x-axis are the zeros of the quadratic function

OpenStudy (moongazer):

If the quadratic function is equal to zero, it will be a quadratic equation

OpenStudy (anonymous):

Where the function touches X axis ( need not cross it)

OpenStudy (anonymous):

is it clear now ???????

OpenStudy (anonymous):

I need 1 answer im getting like 5 different

OpenStudy (anonymous):

if u have to calculate the zeros, then you have to equate the quadratic function = 0 to get the zeros. if you plot a graph, then the point(s) where the graph touches or cuts the x-axis gives the zeros of the function

OpenStudy (moongazer):

i think he needs a one sentence conlusion

OpenStudy (anonymous):

is it clear now??? If you keep looking at answers from various people u will get confused.

OpenStudy (moongazer):

*conclusion

OpenStudy (anonymous):

i still dont get it sorry but what is the answer is it just a point that touches the x axis?

OpenStudy (moongazer):

The graph of a quadratic function is a parabola whose axis of symmetry is parallel to the y-axis. here is a definition from wikipedia but i think it is not a good idea to copy this.

OpenStudy (anonymous):

As I said before, if u r given a graph, then the point(s) where the graph touches or intersects the x-axis, gives the zeros of the quadratic function a Quadratic function can have no real zeros i.e no zero -- In this case the graph will not even touch the x-axis If it has two equal zeros then its tip will just touch the x-axis at one point If it has two distinct zeros then it will cut the x-axis at two places as shown in the file that i attached

OpenStudy (anonymous):

thanks for the medal. for clear picture, attaching a paint file that i made for you, hope it makes my points more clear